OpenStudy (anonymous):

Hi...I seem to have forgotten how to take the derivative of 1/x^2-x+1...can I do a U-sub or is that only for integrating? Need help remembering :o)

OpenStudy (phi):

write it as (x^2-x+1)^ (-1)

OpenStudy (aum):

Is it \(\large \frac{1}{x^2} -x + 1\) or \(\large \frac{1}{x^2-x + 1}\) ?

OpenStudy (phi):

and use d x^n = n x^(n-1) dx

OpenStudy (anonymous):

sorry...I keep losing connection

OpenStudy (anonymous):

thanks everyone...i got it!